### v1

# Does CLIP perceive art the same way we do?

Abstract.
CLIP has emerged as a powerful multimodal model capable of connecting images and text through joint embeddings, but to what extent does it “see” the same way humans do—especially when interpreting artworks? In this paper, we investigate CLIP’s ability to extract high-level semantic and stylistic information from paintings, including both human-created and AI-generated imagery. We evaluate its perception across multiple dimensions—content, scene understanding, artistic style, historical period, and the presence of visual deformations or artifacts. By designing targeted probing tasks and comparing CLIP’s responses to human annotations and expert benchmarks, we explore its alignment with human perceptual and contextual understanding. Our findings reveal both strengths and limitations in CLIP’s visual representations, particularly in relation to aesthetic cues and artistic intent. We further discuss the implications of these insights for using CLIP as a guidance mechanism during generative processes, such as style transfer or prompt-based image synthesis. Our work highlights the need for deeper interpretability in multimodal systems, especially when applied to creative domains where nuance and subjectivity play a central role.
CLIP, multimodal models, painting analysis, generative guidance, vision-language alignment, computational art, visual perception
ccs: Computing methodologies Computer visionccs: Computing methodologies Scene understandingccs: Computing methodologies Neural networksccs: Computing methodologies Learning latent representationsccs: Computing methodologies Visual content-based indexing and retrievalccs: Information systems Multimedia content creationccs: Applied computing Fine arts
Figure 1. 3D UMAP projection of image embeddings of National Gallery of Art Dataset extracted from the CLIP ViT-L/14 model.
[…]
Each point represents a painting from the National Gallery of Art dataset, colored by artistic style. The visualization reveals the semantic organization of the artworks in the latent space, where similar styles tend to cluster together.
1. Introduction
In recent years, multimodal models have reshaped the landscape of machine perception and understanding, with CLIP (Contrastive Language–Image Pretraining) (Radford et al., 2021) standing out as one of the most influential. Trained on hundreds of millions of image–text pairs, CLIP has demonstrated remarkable capabilities across a broad range of tasks, including zero-shot classification (Wang et al., 2025; Li et al., 2025; Yang et al., 2024), image retrieval (Yang et al., 2025; Baldrati et al., 2024; Liu et al., 2024), re-identification (Li et al., 2023; Asperti et al., 2025b), prompt-based generation (Surapaneni and Bhagvati, 2024; Huang et al., 2023), and semantic search (Peng et al., 2024; Zhou et al., 2024). Its success has made it a foundational component in many state-of-the-art generative models and vision-language pipelines. In particular, CLIP’s ability to align visual and textual modalities has been widely adopted as a guiding mechanism in tasks ranging from text-to-image synthesis (e.g., GLIDE (Nichol et al., 2022), DALL·E (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022; Esser et al., 2024)) to creative applications such as style transfer and visual storytelling.
Yet, despite its ubiquity, the nature and limits of CLIP’s perceptual alignment remain underexplored. While CLIP is optimized to match images with descriptive captions, it is less clear whether this alignment genuinely reflects human perception — especially in complex, subjective, or culturally embedded domains like art. Its capacity to identify what is depicted in an image is impressive, but its understanding of how that content is rendered — including factors such as artistic style, visual coherence, or historical context — is far less understood. Moreover, most evaluations have focused on natural images or benchmark datasets, leaving open the question of how well CLIP performs on ambiguous, interpretive, or aesthetic content, such as paintings — particularly those generated by AI.
In this paper, we investigate CLIP’s perceptual capabilities in the domain of visual art, focusing on both human-made and AI-generated paintings. Our goal is to assess how well CLIP captures not only semantic content but also stylistic attributes, temporal signals, and visual deformations. Through a series of probing tasks and analyses, we evaluate CLIP’s representations along multiple interpretive axes, including scene type, artistic style, historical period, and the presence of visual artifacts. To this end, we leverage two richly annotated datasets: classical works from the National Gallery of Art of Washington (of Art, 2024) and synthetic paintings from the AI-Pastiche (Asperti et al., 2025a) collection.
This study serves as a stepping stone toward a broader question: can large vision-language models, such as CLIP, form something akin to an aesthetic sense? Do they internalize representations of style, harmony, or beauty, and if so, are these grounded in visual abstraction, statistical regularities, or biases in their training data?
We offer both a conceptual and empirical investigation into these questions. Our findings reveal a consistent gap between CLIP’s textual associations and its ability to perceive visual nuance. While the model performs well on broad semantic alignment, it struggles with stylistic subtleties, the attribution of artistic periods, and the detection of visual defects in generative imagery. These limitations point to a deeper issue: despite its multimodal power, CLIP lacks a robust internal model of aesthetic form.
By combining metadata-driven evaluation with perceptual baselines, our work presents a structured critique of CLIP’s performance in aesthetic domains. In doing so, we highlight the need for more interpretable and perceptually grounded multimodal systems, particularly as models like CLIP are increasingly used to guide, evaluate, and curate creative content. As these systems become embedded in artistic workflows and cultural interfaces, it becomes crucial to ask not just what they can recognize or generate but how they see — and whether that way of seeing aligns with our own.
2. Related Works
Several recent works have focused on the problem of ascertaining the potential and limitations of CLIP’s vision mechanism and the extent to which the image embedding can be retrieved and effectively exploited through textual prompts.
Concerns about CLIP’s capacity to deeply understand the meaning of its textual prompts were raised in (Pan et al., 2022). In (Abbasi et al., 2025), a comprehensive analysis of CLIP’s performance limitations in multi-object contexts through controlled experiments is presented. The findings reveal significant biases in both encoders: the image encoder favors larger objects, while the text encoder prioritizes objects mentioned first in descriptions.
Several works have been devoted to assess the robustness of CLIP. In (Rosa et al., 2024), the authors investigate the adversarial robustness of CLIP-based methods in detecting AI-generated images, revealing their vulnerability to white-box attacks. A more comprehensive assessment of CLIP’s robustness along different perspectives is done in (Tu et al., 2024): despite several limitations, CLIP models exhibit superior robustness to visual factor-level variations compared to more traditional ImageNet models.
In (Lytvyn et al., 2024), the authors present a CLIP-based method for image similarity evaluation that relies on textual descriptions rather than raw visual features. Using cosine similarity between encoded text and image vectors, the approach captures semantic relationships more effectively than traditional metrics, such as SSIM and FSIM, especially in complex images. The results confirm CLIP’s robustness in tasks such as semantic search and classification, despite some limitations related to model dependence and real-world performance.
A different line of research aimed to investigate the use of CLIP as a guidance technique for image generation. The findings in (Taghipour et al., 2024) suggest that while CLIP embeddings are important for aesthetic quality, they do not contribute significantly to the consistency of the subject and background in the generated outputs. Similar limitations in using CLIP encodings as a synthesis objective are discussed in (Liu et al., 2023), emphasizing the need for a more unified framework for text-guided and image-guided synthesis. None of these articles addresses guidance towards style replication; the conclusion of our work seems to suggest that CLIP is likely of little help for this task, if not altogether detrimental.
The possibility of understanding CLIP internal representations by inverting them with traditional gradient ascent techniques (Mahendran and Vedaldi, 2015; Dosovitskiy and Brox, 2016; Yin et al., 2020) has been explored in (Kazemi et al., 2024). Among other things, the authors remark on the frequent emergence of instances of Not Safe For Work (NSFW) images from the inversion of innocuous prompts. A gradient ascent technique, not based on pixel optimization but on a set of RGBA Bezier curves, is investigated in (Frans et al., 2022).
In light of the previous limitations, several works have been devoted to improving CLIP’s performance, enhancing its downstream generalization ability, and reducing the modality gap between text and images. Adapters technique allow for a lightweight fine-tuning of the model through the insertions of suitable modules. Examples along this direction are CLIP-adapter (Gao et al., 2024), TIP-adapter (Yu et al., 2023), LIxP (Roth et al., 2024) or APE (Zhu et al., 2023). Given that our aim is to explore CLIP’s native perceptual abilities in the context of art, we consider adaptation techniques to be somewhat misaligned with the spirit of our investigation. A more promising line of investigation consists of addressing potential perception issues by focusing on subspaces through suitable projections, as outlined in (Zhu et al., 2024). However, our preliminary investigations in this direction did not yield interesting results.
A different research direction consists of trying to improve the retrieval of multimodal systems through prompt engineering. A common approach followed e.g. in CoOp (Zhou et al., 2022b) or CoCoOp (Zhou et al., 2022a) consists of integrating the tokenization of the prompt with a set of learnable vectors learned by gradient descent. We tested this technique in the case of style classification, without noticeable improvements.
3. Methodology and Data
Our aim is to investigate CLIP’s ability to extract high-level semantic and stylistic information from paintings and to evaluate its perceptual capabilities across multiple dimensions, including content, scene understanding, artistic style, and the presence of visual deformations or artifacts.
We conducted our analyses on two richly annotated datasets: a subset of the freely available National Gallery of Art collection in Washington, and the AI-Pastiche dataset (Asperti et al., 2025a)—a collection of 953 AI-generated paintings produced by 12 different models from 73 carefully crafted prompts, covering a wide range of major artistic styles. Both datasets are described below.
The motivation for comparing human and AI-generated artworks lies in our goal to assess whether CLIP’s vision system can perceive the nuanced differences between them, including the limitations of AI models in accurately replicating artistic styles.
For the paintings from the National Gallery of Art, we leverage the available metadata to evaluate CLIP’s ability to associate each work with a descriptive summary and to understand its artistic style. We employ the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes and Healy, 2018) as a visualization tool for exploring the shared image-text latent space of CLIP. An example of the resulting visualization, where the embeddings of the paintings are color-coded by their respective styles, is shown in Figure 1.
A similar investigation is conducted on the AI-Pastiche dataset, comparing the generated images both to their textual prompts and to the intended artistic styles. This analysis is inherently more complex, as any mismatch between image and prompt could result from either of the two models involved: the image generator, which may fail to accurately follow the prompt, or CLIP, which may fail to correctly identify the intended content or style. To disentangle these factors, we draw upon a set of user surveys conducted by the creators of the AI-Pastiche dataset. These surveys assess the perceived ”authenticity” of the AI-generated artworks, their adherence to the prompt, and the presence of visible artifacts or deformations.
Specifically, we use CLIP to evaluate how closely each generated image aligns with its corresponding prompt and compare these results with human evaluations. Additionally, we test CLIP’s capacity to identify visual defects, such as distortions, inconsistencies, or artifacts commonly produced by generative models.
Most of our investigations were conducted using multiple versions of CLIP, with the secondary goal of comparing their relative performance.
3.1. National Gallery of Art Dataset
The National Gallery of Art Dataset (NGAD) was created by carefully selecting 1,521 artworks from the National Gallery’s publicly available collection, which contains over 130,000 pieces(of Art, 2024). This selection focuses exclusively on paintings and drawings, deliberately excluding sculptures and modern artworks that might be more challenging for models like CLIP to analyze.
Each record in the dataset comprises 11 key attributes, including a unique identifier (objectid), title, creation period (period), artist or attribution, and a high-resolution image link (link) provided via the IIIF protocol. A key feature of the dataset is the description attribute, which provides a detailed textual representation of each artwork. These descriptions are structured to convey visual elements, artistic techniques, and compositional aspects.
The dataset covers a broad range of historical periods and artistic styles. The most represented styles include Baroque (286 artworks), Renaissance (272), and Impressionism (172), reflecting a focus on European artistic traditions.
While key attributes such as objectid, title, period, artist, link, and description are fully populated, certain fields contain missing values. Specifically, keyword is missing in 231 instances, style in 261 instances, and technique in 99 instances.
3.2. AI-Pastiche
The AI-Pastiche dataset (Asperti et al., 2025a) is a curated collection of 953 AI-generated paintings designed to mimic historical artistic styles. Created using 73 meticulously crafted prompts, the dataset spans a wide range of art periods. The images were generated using multiple state-of-the-art generative models and are intended to support the evaluation of how convincingly AI can reproduce the visual characteristics of human art. Each image is accompanied by rich metadata, including the generative model used, the prompt text, intended style and period, as well as a list of subject descriptors like ”crowd”, ”landscape” or ”soft tones.” The dataset enables research on stylistic imitation, model benchmarking, and user perception studies. It serves as a valuable resource for analyzing the capabilities and limitations of generative models in the artistic domain and is publicly available to support further exploration in AI-driven art.
4. Experiments on human artworks
The experiments in this section were conducted on a subset of the National Gallery of Art (Washington). We designed two distinct experiments. The first aims to assess CLIP’s ability to associate each image with a summary of its corresponding description. The second focuses on evaluating CLIP’s capacity to distinguish between different artistic styles.
4.1. Image-description alignment
In the first experiment, we encountered a limitation: CLIP accepts a maximum of 77 tokens as input, whereas the painting descriptions from the National Gallery of Art often exceeded this limit. To address this, we generated concise summaries of each description using ChatGPT 4o-mini. We passed as input to ChaptGPT also the subject, the style and the period of the painting, asking him to retain this information in the summary. The actual prompt used to generate the image is discussed in the appendix.
For each image-summary pair $\langle x,s\rangle$, we computed the cosine similarity between $CLIP_{image}(x)$ and $CLIP_{text}(s)$ and used this value as a ranking score to compute recall at different thresholds.
The experiments were repeated for several different versions of CLIP available in the OpenAI library(Radford et al., 2021).
The results of this experiment are shown in Table 1.
Model recall@1 recall@5 recall@10
RN50 0.663 0.915 0.966
RN101 0.693 0.926 0.966
RN50x4 0.741 0.946 0.978
RN50x16 0.791 0.964 0.988
RN50x64 0.828 0.97 0.99
ViT-B/32 0.678 0.925 0.97
ViT-B/16 0.709 0.928 0.969
ViT-L/14 0.794 0.972 0.989
ViT-L/14@336px 0.814 0.974 0.991
Table 1. Summary-image alignment for NGAD images.
These results show a consistent improvement in performance as the capacity of the CLIP models increases. Among the ResNet-based architectures, recall metrics progressively increase from RN50 up to RN50x64, with RN50x64 achieving the best performance in this family. Similarly, for the ViT-based models, the larger architectures and higher input resolutions result in better alignment, with ViT-L/14@336px showing the highest performance among the transformer variants.
Overall, while the ResNet-based RN50x64 yields the best recall@1 scores in this experiment, ViT-L/14@336px performs comparably, especially at higher recall thresholds. These results suggest that both architectural complexity and input resolution play a crucial role in enhancing the image-text alignment capabilities of CLIP, particularly in tasks involving fine-grained associations, such as matching painting summaries with artworks.
4.2. Style Recognition
In this second experiment, we assess CLIP’s ability to associate artworks with their corresponding artistic styles. For each unique style present in the dataset, we generated a fixed textual prompt in the form “an artwork in [style] style”. Using CLIP’s image and text encoders, we computed normalized embeddings for both modalities. The cosine similarity between text and image embeddings was then calculated to produce a similarity matrix, capturing the degree of alignment between each image and every style prompt.
Performance was evaluated using recall@$k$ metrics, which measure the proportion of test images for which the correct style appears among the top-$k$ most similar text prompts. Specifically, recall@$k$ is defined as the percentage of images for which the ground-truth style is ranked within the top $k$ most similar entries in the corresponding column of the similarity matrix.
Table 2 reports the results obtained across a range of CLIP architectures. Among all tested models, ViT-L/14@336px achieved the best performance, with a recall@1 of 0.457 and a recall@5 of 0.831. However, the results across all models remain moderate, highlighting the increased complexity of style recognition compared to textual description alignment.
model recall@1 recall@2 recall@3 recall@5
RN50 0.354 0.546 0.662 0.801
RN101 0.379 0.577 0.674 0.78
RN50x4 0.344 0.504 0.611 0.772
RN50x16 0.373 0.581 0.679 0.786
RN50x64 0.343 0.516 0.627 0.766
ViT-B/32 0.316 0.467 0.585 0.737
ViT-B/16 0.349 0.506 0.622 0.766
ViT-L/14 0.4 0.577 0.697 0.795
ViT-L/14@336px 0.457 0.632 0.716 0.831
Table 2. Recall@k scores for art style recognition on NGAD.
To better understand the nature of the model’s errors, we present in Figure 2 the normalized confusion matrix obtained from the best-performing model. While the overall recall scores may appear modest, a closer inspection of the confusion matrix provides important nuance. Notably, many misclassifications occur between adjacent or stylistically related categories. For instance, Impressionism is frequently confused with Post-Impressionism, Abstract Expressionism with Minimalism, and Baroque with Rococo.
Interestingly, the Academic style appears to serve as a kind of fallback category, absorbing a diverse set of misclassifications from styles such as Baroque, Neoclassicism, Romanticism, and Realism. This suggests that the model may associate ”academic” with a broad range of classical or representational visual features, especially when more specific stylistic signals are ambiguous or absent.
Figure 2. Normalized confusion matrix of the best CLIP model (ViT-L/14@336px) on style classification.
This can be better observed in Figure 3, where we compare the distribution of true and predicted styles across the dataset. This analysis highlights notable disparities, reflecting both the imbalance present in the dataset and the most frequent mistakes in the model’s predictions.
Figure 3. Comparison between actual and predicted distribution of styles in NGAD.
These findings suggest that while CLIP demonstrates a degree of sensitivity to artistic style, its current representations are not fully suited for tasks requiring a nuanced understanding of artistic conventions or visual grammar, warranting further refinement or supervision for such objectives.
A qualitative inspection of misclassified examples offers further insight into CLIP’s limitations in style recognition. Figure 4 showcases three representative failure cases. Each image is shown with its true style label and the incorrect prediction made by the best-performing model.
Refer to caption Refer to caption Refer to caption
True: Impressionist Pred: Renaissance True: Realist Pred: Rococo True: Orientalist Pred: Neoclassic
Figure 4. Examples of misclassifications in style recognition
These examples illustrate how overlapping visual features or contextual ambiguity can lead to misclassifications. In several cases, the predicted style may appear visually plausible, emphasizing the challenge of the task.
4.3. Semantic Relationships in the Latent Space
To investigate the internal organization of the latent space learned by CLIP, we employed the UMAP algorithm (McInnes and Healy, 2018) to generate a three-dimensional projection of the high-dimensional embeddings. These embeddings included both image representations and textual prompts describing artistic styles (e.g., ”an artwork in [style] style”). The resulting visualization (Figure  5) reveals a clear separation between textual encodings (the small cluster on the left) and image encodings (the large cluster on the right). For the images, we also distinguish correctly classified samples, shown as green bullets, from misclassified ones, shown as red crosses.
The substantial entanglement of the two image classes suggests a dominance of non-stylistic features in the embeddings. However, the chaotic pattern could also be a consequence of the aggressive dimensionality reduction and may not accurately reflect the semantic structure present in the original high-dimensional space.
Figure 5. Three-dimensional projection of the textual embeddings of artistic styles and the visual embeddings of correctly classified (green) and misclassified (red) images using UMAP.
To gain a deeper understanding of the structure of CLIP’s latent space, we conducted a detailed analysis based on nearest neighbors. Specifically, we focused on image pairs in which one image was correctly classified while the other was not. By identifying such pairs, we were able to isolate semantically coherent examples where classification performance diverged. This setup offered valuable insight into how subject matter similarity and stylistic cues interact within CLIP’s representation space, and where the model may struggle to disentangle the two.
Representative examples are shown in Figure 6, where the misclassified image is displayed on the left, and its nearest correctly classified neighbor appears on the right. For each image, we report the cosine similarity scores with respect to both the true and the predicted style prompts, computed in both the original latent space and its lower-dimensional projection. This comparison allows us to examine how proximity in the embedding space relates to classification outcomes and whether stylistic distinctions are preserved through dimensionality reduction.
Figure 6. Visual comparison between a misclassified image (left) and a correctly classified image (right). For each artwork, the actual and predicted artistic styles are shown, along with their respective similarity scores to the image in the latent CLIP space.
These findings underscore a key limitation of CLIP: it tends to encode and prioritize semantic content—like objects, scenes, and compositions—over stylistic features such as brushwork, color palette, or compositional structure. This leads to frequent misclassifications when artworks differ in style but share similar subject matter. This bias is consistent with CLIP’s training objectives, which favor semantic alignment based on large-scale image-text data, where captions often emphasize content over formal style. As a result, CLIP’s latent space tends to conflate stylistically diverse images with similar semantics. While we explored projections to mitigate this issue, the results were unsatisfactory. Techniques like light adapters could potentially fine-tune CLIP’s vision encoder; such modifications fall outside the scope of our current research aims.
5. Experiments with AI-generated artworks
The experiments conducted on AI-generated images follow a structure similar to those applied to human-generated artworks, but with some important caveats. In this case, the prompt used to generate each image serves as the reference summary for computing similarity. Consequently, a low similarity score may indicate a failure on the part of the image generator rather than a shortcoming of CLIP.
The same ambiguity arises in style classification: in the AI-Pastiche dataset, the ”style” corresponds to the intended style described in the prompt—not necessarily the one successfully rendered in the generated image. As such, any misclassification could be due either to the image generator failing to follow the prompt or to CLIP failing to recognize the intended style.
To clarify the contribution of these factors, we further compared CLIP’s perception of generated images with human judgments, drawing on user survey data from the AI-Pastiche dataset (Asperti et al., 2025a). Section 6 provides details of these experiments.
5.1. Image-prompt similarity
Here, we are comparing the embedding of the generated image with the embedding of the relative prompt. AI-pastiche prompts are quite articulated, and we faced a similar problem to the case of real image descriptions due to the maximum numbers of tokens accepted by CLIP. For this reason, we created a prompt-summary in a way similar to the technique described in Section 4.1.
We generated embeddings for both images and prompt summaries and computed their cosine similarity, using this value to predict the prompt for each generated image. Due to the smaller number of textual prompts (72), we just measure accuracy in this case.
The results are shown in Table 3.
Model RN50 RN101 RN50x4 RN50x16 RN50x64 ViT-B/32 ViT-B/16 ViT-L/14 ViT-L/14@336px
Accuracy 0.866 0.887 0.891 0.893 0.896 0.881 0.880 0.896 0.896
Table 3. Accuracy of different CLIP models in matching generated images with their corresponding summarized prompts in the AI-Pastiche dataset.
Model RN50 RN101 RN50x4 RN50x16 RN50x64 ViT-B/32 ViT-B/16 ViT-L/14 ViT-L/14@336px
Accuracy 0.467 0.455 0.458 0.448 0.376 0.443 0.437 0.470 0.487
Table 4. Result table for art style recognition on AI-Pastiche
All CLIP models perform well in associating generated images with their corresponding prompt summaries, with accuracy values exceeding 0.86 across the board. The highest performance is achieved by RN50x64, ViT-L/14, and ViT-L/14@336px, all reaching an accuracy of 0.896. The task is sensibly simpler than the image-description alignment of Section 4.1, since we only have 73 prompts relative to quite different subjects. Nevertheless, the results confirm CLIP’s ability to capture the visual-semantic correspondence in synthetically generated image-text pairs.
From the perspective of the generators, the high recall indicates that all models performed well in producing images that closely matched the subjects described in the prompts. The average cosine similarity between each generated image and its corresponding prompt summary is 0.278, with a standard deviation of 0.344.
Although cosine similarity can be profitably used to associate an image to its prompt, it is not clear if it can be reliably employed as a standalone metric to assess the quality of different generated images in terms of their alignment to the same prompt. The problem is that this assessment requires a complex evaluation comprising not just the semantic correspondence with the subject but also the stylistic adherence and the technical quality of the generation. This includes evaluating the absence of artifacts, distortions, or visual defects that may not be compatible with the intended artistic style.
We start addressing stylistic issues in Section 5.2; and in Section 6.1 we will compare the CLIP-evaluation of the adherence between an image and its prompt with a similar evaluation done by human experts.
5.2. Style Recognition
In the second experiment, we used CLIP to evaluate the alignment between generated artworks and their expected styles, provided among the AI-Pastiche metadata. Similar to the case of images from NGAD, we generated a prompt of the form ”an artwork in [style] style” and computed its cosine similarity with the image embedding, deriving a confusion matrix. The accuracy results for the different models are shown in Table 2.
Considering that we are comparing the predicted style of generated images with the target style specified in the prompt, the classification accuracy achieved by CLIP is surprisingly high and comparable to the accuracy observed in the NGAD task.
Upon visual inspection, the generators evaluated in the AI-Pastiche dataset do not appear to reproduce the historical styles specified in the prompts convincingly.
In this context, the high agreement between CLIP’s classification and the original prompt may reflect a shared inductive bias rather than genuine stylistic fidelity: the reason is that many generative models rely on CLIP during training or inference as a scoring function, conditioning mechanism, or similarity guide.
In analogy with the work done for NGAD, a confusion matrix was constructed to better analyze CLIP model for predicting style in AI-Pastiche (Figure 7) and the distribution histograms of the actual and predicted values for AI-Pastiche (Figure 8).
Figure 7. Confusion matrix of the best CLIP model for predicting style in AI-Pastiche
Figure 8. Distribution of actual values and predicted values in AI-Pastiche
The distribution of actual and predicted styles in AI-Pastiche (Figure 8) reveals acceptable performance on several of the most frequently prompted styles, such as Renaissance, Impressionism, Surrealism, and Cubism, while styles like Romanticism, Dadaism, and Classicism are more often misclassified.
To better understand how CLIP assessments compare with human perception, we turn to the adherence and artifact surveys provided in the AI-Pastiche dataset. These human evaluations offer a valuable reference point for gauging whether CLIP captures stylistic and visual cues in a way that aligns with human judgment.
6. Comparison with human evaluations
In this section, we directly compare CLIP’s perceptual judgments with human interpretations, drawing on the results of two user surveys from the AI-Pastiche dataset: the adherence survey, which measures how closely generated images align with their prompt, and the artifact survey, which evaluates the presence of visual distortions or inconsistencies. These comparisons provide a human-centered benchmark for evaluating CLIP’s ability to detect stylistic cues and perceptual artifacts in generated artwork.
6.1. Adherence Analysis
In (Asperti et al., 2025a), an adherence survey was conducted in which participants were shown a set of images generated by different models from the same textual prompt. They were asked to rate each image as good, neutral, or bad in terms of its adherence to the prompt in a comparative setting. By averaging responses from multiple participants, the authors derived an adherence score for each image, reflecting perceived stylistic and semantic alignment with the original prompt.
To explore whether CLIP captures similar perceptual cues, we conducted a parallel analysis using CLIP similarity scores. For each prompt, we computed the cosine similarity between the CLIP text embedding of the prompt and the CLIP image embeddings of all associated generated images. These scores were then normalized and scaled to produce a vector directly comparable to the human-derived adherence scores.
Finally, we computed the correlation between the CLIP similarity vector and the human adherence scores, using it as a synthetic measure of alignment between CLIP’s assessments and human perception. This procedure was repeated across all available CLIP models, and the results are reported in the second column of Table 5 (the third column is discussed in Section 6.2).
Model Correlation with human evaluation Integration of defects
RN50 0.406 0.478
RN101 0.398 0.462
RN50x4 0.379 0.448
RN50x16 0.413 0.489
RN50x64 0.428 0.484
ViT-B/32 0.411 0.481
ViT-B/16 0.383 0.458
ViT-L/14 0.425 0.482
ViT-L/14@336px 0.437 0.497
Table 5. Correlation between a human evaluation of the adherence between a generated image and its prompt, compared with a similar evaluation based on CLIP’s.
Across all models, the similarity scores are around 0.4. The result is not too low, but it should be compared with the correlation between evaluations of different humans, which on average is around 0.7, which is sensibly higher.
6.2. Perception of Artifacts and Deformations
According to the instruction provided in (Asperti et al., 2025a), the user was asked to perform an adherence evaluation essentially driven by three different directions:
  1. (1)
    Content: respect of the subject described in the prompt:
  2. (2)
    Style: respect of the style that the generative model was required to imitate;
  3. (3)
    Overall Quality: absence of visible artifacts and deformations typical of generative techniques.
One of the surveys collected in (Asperti et al., 2025a) was specifically aimed at detecting the presence of visible artifacts or deformations in the generated image. Defects were categorized as Major (clearly visible of frequent errors, such as macroscopic anatomical mistakes), Minor (additional fingers, minor deformations), or None (no apparent mistake). Results were summarized in a ”defect score” associated with each image.
We sought to investigate whether CLIP could identify and detect these kinds of mistakes. Our initial investigation aimed to determine if we could approximate the defect score through linear regression, starting from the CLIP embedding of the AI-Pastiche images.
The result was negative: we obtained a coefficient of determination $R^{2}$ close to 0.
As additional evidence that CLIP embeddings do not account for defects and artifacts in input images, we test whether CLIP’s evaluation of prompt adherence can be improved by incorporating a linear combination with the human-evaluated ”defect score”. This turns out to be the case: a suitable linear combination achieves a similarity of nearly 0.5 with the average human adherence evaluation (see the third column in Table 5).
Nevertheless, CLIP sometimes assigns high adherence scores to images that humans evaluate poorly and vice versa. This discrepancy is often due to mismatches in content or style. Some illustrative examples are shown in Figure 9.
Refer to caption Refer to caption Refer to caption
(a) Midjourney (b) Auto-Aesthetics (c) Omnigen
Figure 9. Examples of images in the AI AI-Pastiche dataset not aligning well with their prompts, for content or style.
Figure (a) was supposed to represent a fight between two knights on horseback; humans penalized the fact that only one knight is represented, while CLIP does not seem to notice the absence. Image (b), generated by Auto-Aesthetics V1, was meant to depict a lively Parisian street scene in the rain, in a style resembling impressionism: humans reproach the lack of adherence to the required style. The prompt for image (c) required the representation of a kneeling figure in rich robes offering a white flower to another figure, draped in blue; the fact that one of the two figures is missing is not captured by CLIP similarity.
7. Conclusion
Our investigation into CLIP’s perception of artworks—across both human-created and AI-generated images—reveals a model with remarkable breadth, but one still far from capturing the richness of human aesthetic and contextual understanding. While CLIP is adept at grounding images in broad semantic categories and descriptive summaries, it often falters when asked to navigate the more subjective terrain of artistic nuance: style, intention, emotional tone, and cultural context.
This gap is particularly evident in generative contexts, where CLIP must assess not only what is depicted, but how and why. When the prompt serves as both the generative instruction and the evaluative anchor, the act of measuring alignment becomes ambiguous: are mismatches due to the generator’s limitations or CLIP’s perceptual blind spots? Our experiments, informed by human judgments, suggest that CLIP often lacks the sensitivity required to discern these subtle failures—especially when artifacts are stylistically ”plausible” but semantically off-key.
More fundamentally, our work raises questions about how we interrogate multimodal representations. CLIP’s joint embedding space promises a bridge between vision and language, yet that bridge is asymmetrical: text embeddings are inherently structured and interpretable, while image embeddings remain opaque and spatially abstract. Probing image understanding through text similarity metrics can therefore be misleading, especially when the image conveys signals that elude linguistic capture—such as texture, composition, or emotional affect. In the case of art, where meaning is often non-verbal and deliberately ambiguous, this mismatch becomes a critical limitation.
Looking ahead, we believe that future vision-language systems must go beyond mere alignment. What is needed is a deeper model of perception—one that can reason about images not just in terms of objects or styles, but in terms of historical context, artistic intent, and visual storytelling. This may require models trained with richer supervision, involving not only captioned data but also art historical metadata, expert narratives, and multimodal dialogues.
Ultimately, as AI is increasingly used to create, curate, and critique visual culture, we must ask not just what models see but how they see, and whose eyes they are borrowing. CLIP represents a powerful beginning, but it is not yet a substitute for human perception in the arts. Rather, it is a lens—partial, biased, but illuminating—that can help us better understand both machine vision and our own ways of seeing.
Acknowledgements.
This Research has been partially supported by the Future AI Research (FAIR) project of the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 funded from the European Union - NextGenerationEU.
8. Appendices
Appendix A Example of an artwork from the National Gallery
To illustrate the methodology adopted in our study, we selected a painting from the National Gallery of Art, Van Gogh’s self-portrait. Since zero-shot CLIP accepts textual descriptions of no more than 77 tokens, we employed GPT-4o mini to generate a summary of the description of the painting, restricted to a maximum of 300 characters. The summarization process aimed to preserve essential information related to the subject and the stylistic characteristics of the artwork. This is the actual prompt used to generate the summaries:
Your goal is to summarize the following painting descriptions in 300 characters.
You will be provided the description of a painting, its subject, its style, and its period, and you will output a JSON object containing the following information:
summary: string // at most 300 characters summary of the painting based on the painting description.
The summary must retain information about the subject, style, and period.
In Figure  10 we present the painting, the official description provided by the National Gallery, and the summarized version produced using GPT-4o mini.
Figure 10. Example of an artwork from the National Gallery.
National Gallery Description:
Shown from the chest up, a man with short, orange hair and green-tinted, pale skin looks at us, wearing a vivid blue painter’s smock in this vertical portrait painting. His smock and the background are painted with long, mostly parallel strokes of cobalt, azure, and lapis blue. His shoulders are angled to our left, and he looks at us from the corners of his blue eyes. He has a long, slightly bumped nose, and his lips are closed within a full, rust-orange beard. He holds a palette and paintbrushes in his left hand, in the lower left corner of the canvas. The background is painted with long brushstrokes that follow the contours of his head and torso to create an aura-like effect.
Summarized version of the description:
This Post-Impressionist portrait by Vincent van Gogh (1876-1900) depicts a man with orange hair and pale skin in a blue painter’s smock. He gazes at the viewer with blue eyes, holding a palette and brushes. The background features long, parallel strokes of blue, enhancing the aura around him.
Appendix B Example of an AI-generated artwork from the AI-Pastiche Dataset
As an example of the AI-Pastiche dataset, we chose an Impressionist painting generated by generative models, depicting a bridge landscape. As observed in the artwork, the generative model adhered to the given prompt, producing an image strongly reminiscent of Monet’s Japanese bridge. For the CLIP processing, the same procedure was applied to the description, limiting it to 300 characters of text, as done for the National Gallery descriptions.
In Figure  11, we present the generated painting, the prompt provided to the model, and the corresponding summary.
Figure 11. Example of an artwork from the AI-Pastiche.
Prompt for the generated image:
Create an Impressionist-style painting depicting a serene outdoor scene, such as a sunlit garden, a riverside, or a city park. The image should focus on capturing the play of natural light and atmosphere, with soft, loose brushstrokes and a pastel-like color palette of light blues, greens, pinks, and yellows. The figures and landscape should appear slightly blurred, as if seen from a distance, giving a sense of movement and fleeting moments. Include reflections in water, dappled sunlight, and subtle shifts in color to evoke a peaceful, idyllic mood, typical of Impressionist art.
Summarized version of the prompt:
This Impressionist painting from the XIX century features a serene landscape with soft tones, depicting a sunlit garden or riverside. It captures natural light and atmosphere through loose brushstrokes and a pastel palette. Figures and scenery appear slightly blurred, evoking movement and tranquility, with reflections in water.
Appendix C The distribution of labels predicted by all CLIP models
Figure 12 presents the style prediction results for all CLIP models. As shown in the graph, most models exhibit challenges in accurately identifying Romantic, Art Nouveau, and Dadaist styles. Notably, many models overestimated the prevalence of Romantic artworks compared to their actual representation in the dataset. The underperformance in predicting Art Nouveau and Dadaist styles may stem from their visual similarities to avant-garde movements such as Futurism, Abstract art, Cubism, and Surrealism, which could lead to misclassifications. However, as observed in our analysis, the ViT 14@336px model demonstrates the strongest overall performance, achieving higher accuracy and robustness across style categories compared to other architectures.
Figure 12. Histogram of style prediction of all CLIP models on AI-Pastiche
Appendix D CLIP Image embedding visualization
Figure 13. 3D UMAP projection of image embeddings of National Gallery of Art Dataset extracted from the CLIP ViT-L/14 model.
Each point represents a painting from the National Gallery dataset, colored by artistic style. The visualization reveals the semantic organization of the artworks in the latent space, where similar styles tend to cluster together.
Appendix E Example of a wrong summary associated with an artwork of the National Gallery of Art
Figure 14. Example of a misclassified summary
In Figure 14, there is an example of misclassifications in the summary recognition of an artwork in the National Gallery Dataset of Art computed using the CLIP ViT-L/14 model.
Actual summary:
In this Baroque painting by Canaletto (1751-1775), sunlight illuminates a vibrant landscape featuring an arched stone ruin, a bridge, and a river. Scattered figures in colorful attire engage with nature, while buildings and churches rise on a distant hill. The scene is rich with detail, showcasing Venetian life and architecture.
Predicted summary:
In this Rococo painting by Jean Honoré Fragonard (1751-1775), a lively scene unfolds in a lush park where light-skinned figures enjoy leisure by a river. A couple in elegant attire sits nearby, while boys engage in playful horse-riding games. Tall trees frame the idyllic landscape, enhancing the theme of amusement.
Appendix F Example of artifact presence in AI-generated artwork from the AI-Pastiche dataset
Figure 15 displays an AI-generated painting depicting Madonna and Child. As is characteristic of synthetic artworks produced by current generative models, the image exhibits pronounced anatomical distortions, particularly in the rendering of hands and facial features. A key objective of our study is to investigate whether CLIP-based analysis can directly extract meaningful information about such artifacts from image embeddings, bypassing manual annotation. By probing the latent representations of these synthetic outputs, we aim to identify patterns associated with structural anomalies (e.g., incoherent geometries and asymmetries) and evaluate their utility for automated quality assessment or iterative refinement of generative models.
Figure 15. AI-generated artwork from the AI-Pastiche dataset depicting Madonna and Child, showcasing anatomical artifacts.